If C=700+0.8Yd, I= 400, Tax=10%, & G= 200, then:

Derive equation for AD
Determine equilibrium value of output.
Calculate value of multiplier
Repeat first 3 parts if Govt. spending is increased to 300
Show changes through graph..
1
Expert's answer
2020-05-11T19:36:41-0400

1)

"AD=C+I+G"

"Yd=Y-T*Y"

"AD=700+0.8(Y-0.1Y)+400+200=1300+0.72Y"

2)

"AD=Y"

"Y=1300+0.72Y"

"Y-0.72Y=1300"

"0.28Y=1300"

"Y=1300\/0.28=4642.86"

3)

M-value of multiplier 

"M=1\/0.28=3.57"

4)

"AD=1300+300+0.72Y=1600+0.72Y"

"Y=1600+0.72Y"

"Y=1600\/0.28=5714.29"

"M=1\/0.28=3.57"

5)